The population of a village was 8000 at the beginning of 2015. If 1600 people were migrated here from our place at the beginning of 2016, find the population at the end of  2017 if the population growth rate was 5% yearly.

Grade:10

2 Answers

Arun
4 years ago
Dear student
 
in the beginning of 2016 = 8000 + 8000*5/100 = 8400
Now 1600 people migrated, hence remaining = 8400 – 1600 = 6800
now at the end of 2017 = 6800 + 6800*5/100 = 7140
Vikas TU
4 years ago
Population in the year 2015 = 8000
Population rate growth = 5% 
So, in 2016 popultion should be 8000+ 8000*5/100 = 8400
Remaining people after migration = 8400-1600 = 6800 
5% populaton will increase at the end of 2017 
6800+ 6800*5/100 = 7140 would be correct ans